Rectangle has a width of 5 and height of x, if the width of the rectangle is doubled and its length is halved, thent he area os the new rectangle is given by what formula?
Answer by alicealc(293)

Width = 5
Length = x

Width is doubled -> new width = 2*5 = 10
Length is halved -> new length = 1/2 * x = x/2

new area = new length * new width = x/2 * 10 = 10x/2 = 5x